OpenStudy (anonymous):

A 5'6" person walking down the street notices his shadow. If the angle of elevation from the tip of the shadow to the sun is 60°, what is the length of the shadow(round to 2 decimal places)? A) 3.18 feet B) 3.23 feet C) 8.66 feet D) 11.00 feet

OpenStudy (mertsj):



OpenStudy (mertsj):

I see that you are studying the special right triangles. You should know that the side opposite the 30 degree angle can be multiplied by sqrt3 to get the long leg. Therefore take the long leg and divide by sqrt3 to get the short leg.

jimthompson5910 (jim_thompson5910):

5' 6" = 5 feet + 6 inches = 12*5 + 6 = 60 + 6 = 66 inches tan(60) = 66/x x*tan(60) = 66 x = 66/tan(60) x = 66/sqrt(3) x = 66*sqrt(3)/3 x = 38.105 So the length of the shadow is roughly 38.105 inches Convert to feet to get 38.105/12 = 3.175 feet which rounds to 3.18 feet

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

im still trying to work it out

OpenStudy (mertsj):

Just take 5.5 divided by square root of 3

OpenStudy (anonymous):

B) 3.23 feet

OpenStudy (anonymous):

i still dont really get it though

OpenStudy (anonymous):

and this is a retake for extra credit